OpenStudy (anonymous):

The price of a mobile phone decreased from $375 to $250 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 32% B. 33% C. 50% D. 83%

OpenStudy (anonymous):

You can solve this with a proportion: \[\frac{ change }{ original }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

125/375 = x/100

OpenStudy (anonymous):

33%